### v1

# Fair Mixed Effects Support Vector Machine

Abstract.
To ensure unbiased and ethical automated predictions, fairness must be a core principle in machine learning applications. Fairness in machine learning aims to mitigate biases present in the training data and model imperfections that could lead to discriminatory outcomes. This is achieved by preventing the model from making decisions based on sensitive characteristics like ethnicity or sexual orientation. A fundamental assumption in machine learning is the independence of observations. However, this assumption often does not hold true for data describing social phenomena, where data points are often clustered based. Hence, if the machine learning models do not account for the cluster correlations, the results may be biased. Especially high is the bias in cases where the cluster assignment is correlated to the variable of interest. We present a fair mixed effects support vector machine algorithm that can handle both problems simultaneously. With a reproducible simulation study we demonstrate the impact of clustered data on the quality of fair machine learning predictions.
Key words and phrases:
Support Vector Machine, Fair Machine Learning, Mixed Models.
2020 Mathematics Subject Classification:
90C90, 90-08, 68T99
1. Introduction
The rise of automated decision-making systems calls for the development of fair algorithms. These algorithms must be constrained by societal values, particularly to avoid discrimination against any population group. While machine learning offers efficiency gains, it can also unintentionally perpetuate biases in critical areas like loan approvals ([das2021fairness]) and criminal justice ([green2018fair]). In loan applications, factors like marital status can lead to unfair disadvantages for single individuals, while in criminal justice, algorithms might associate race with recidivism risk, leading to discriminatory sentencing despite individual circumstances. This highlights the need for fair and unbiased AI frameworks to ensure equal opportunities and outcomes for all.
[…]
Driven by the need to mitigate bias in algorithms, fair machine learning is experiencing a surge in research activity. Numerous research articles are exploring innovative approaches to achieve fairer outcomes. Notable examples include fair versions of Logistic and Linear Regression ([fairRL2]), Support Vector Machines ([FairSVM2]), Random Forests ([fairRF]), Decision Trees ([FairDT]), and Generalized Linear Models (GLMs) ([fairGLM]). These methods aim to address potential discrimination arising from historical data or algorithmic design, ensuring fairer outcomes for all individuals.
In this paper we propose a Mixed Effects Support Vector Machine for fair predictions. We show how to estimate the model and evaluate its performance against the current model that does not take the possible clustering of the data into account. Similar approaches for longitudinal data can be found in [SVM1, SVM3], for Least-squares support vector machine in [SVM5] and applications in agricultural activities in [SVM2].
The paper is organized as follows: In Section 2 we will explore the theory and metrics behind fairness in machine learning and how it applies to support vector machines. In Section 3 we establish the theoretical underpinnings of fair mixed effects support vector machine and propose a strategy for solving them. In Section 4, we conduct a comprehensive evaluation of our proposed method’s effectiveness through various tests. Finally, in Section 5, we demonstrate the practical applicability of our algorithm by solving a real-world problem using the Adult dataset ([Adult]). Our key findings and potential future directions are presented in Section 6.
2. Fair Support Vector Machine
In a binary classification, identifying a function that accurately predicts a label $\hat{y}\in\{0,1\}$ from a feature vector $x\in\mathbb{R}^{p}$ is paramount. This objective is achieved by training the model on a dataset $D={(x^{\ell},y_{\ell})}^{N}_{\ell=1}$.
In the realm of fair binary classification, each observation $x^{\ell}$ possesses a corresponding sensitive attribute $s_{\ell}$ a binary value of either 00 or $1$. The goal becomes identifying a solution that balances accuracy with fairness. Fairness in machine learning can be assessed using various metrics, but this discussion focuses on disparate impact (DI). Alternative fairness metrics for binary classifiers can be found in [Paper1].
A binary classifier is deemed to be free from disparate impact if the probability of it classifying an instance as either 00 or $1$ is identical for both values of the sensitive feature being considered. In other terms, disparate impact is absent when the classifier treats all groups equally regardless of their sensitive feature values. Mathematically, we can write this as:
$P(\hat{y}_{\ell}=1|s_{\ell}=0)=P(\hat{y}_{\ell}=1|s_{\ell}=1).$
An algorithm exhibits disparate impact when it produces significantly more favorable outcomes, for certain groups defined by sensitive features. In the following we present how to add the fairness in the context of Support Vector Machine (SVM).
In SVM, proposed by [vapnik1964class] and [hearst1998support], a hyperplane $\beta^{\top}x$ splits the data based on its classification. Note that in our formulation the first column of X is equal to 1 everywhere and, hence, the bias parameter is the first entry in $\beta$ ([hsieh2008dual]). By adding the fairness constraint, our fixed effect $\beta$ can be found by solving the following quadratic optimization problem.
$\underset{\left(\beta,\xi\right)}{\min}$ $\;\;\;\;\frac{1}{2}\|\beta\|^{2}+K\sum_{\ell=1}^{N}\xi_{\ell}$ (1)
$\rm{s.t}$ $\;\;\;\;y_{\ell}(\beta^{\top}x_{\ell})\geq 1-\xi_{\ell},\text{ }\ell=1,\dots,N,$ (2)
$\;\;\;\;\frac{1}{N}\sum\limits_{\ell=1}^{N}(s_{\ell}-\bar{s})(\beta^{\top}x_{\ell})\leq c,$ (3)
$\;\;\;\;\frac{1}{N}\sum\limits_{\ell=1}^{N}(s_{\ell}-\bar{s})(\beta^{\top}x_{\ell})\geq-c,$ (4)
$\;\;\;\;\xi_{\ell}\geq 0,\text{ }\ell=1,\dots,N,$ (5)
with
$\bar{s}=\dfrac{\displaystyle\sum_{\ell=1}^{N}s_{\ell}}{N},$ (6)
and $s$ being the sensitive feature. Constraints (1), (2) and (5) are from the Support Vector Machine problem. Constraints (3) and (4) guarantees the fairness in the prediction. The construction and justification of these constraints can be found in [pmlr-v54-zafar17a]. The parameter $c\in\mathbb{R}^{+}$ is a threshold that controls the importance of fairness. We refer to the problem above as Fair Support Vector Machine (SVMF).
3. Fair Mixed Effects Support Vector Machine
Often, data exhibits unexplained variability within groups or at different levels. While these “random effects” remain unobserved directly, statistical methods can estimate their impact. They represent deviations from the overall trend specific to individuals or groups as we can see in clustered or stratified data. To account for this variability, we include them into our optimization problem. For that, we incorporate both fixed effects (average effect) and random effects in our SVM model.
First we need incorporate, into the Objective function (1), the random effects $g_{i}$, with $i$ being the random effect group, where $i\leavevmode\nobreak\ \in\leavevmode\nobreak\ [1,n]\leavevmode\nobreak\ :=\leavevmode\nobreak\ \{1,\dots,n\}$ and $n$ the number of groups. We can add random effects to the objective function by adding a term that penalizes the variance of the random effects. This helps to ensure that the parameters of the model are estimated in a way that minimizes bias and provides accurate and reliable results. Since the variance of $g$ is:
$\sigma^{2}=\frac{\sum_{i=1}^{n}(g_{i}-\bar{g})^{2}}{n-1},$
and the random effect $g$ follows a normal distribution with mean zero, then:
$\sigma^{2}$ $=\frac{\sum_{i=1}^{n}(g_{i}-0)^{2}}{n-1}=\frac{\sum_{i=1}^{n}g_{i}^{2}}{n-1}.$
This means that our regularization term is $\frac{\sum_{i=1}^{n}g_{i}^{2}}{n-1}$. Since we have a minimization problem and $n-1$ is fixed, we can consider w.l.o.g a parameter $\lambda$ that controls the importance of the random effects.
In the Constraint (2) and in the fairness constraints (3) and (4), we also add the random effect. This allows the fixed effects $\beta$’s to find the hyperplane that separates the data, while the random effects allow this hyperplane to vary in position depending on the group. So each group has the same fixed effect but different random effect.
Consequently, we formulate the following constrained optimization problem:
$\begin{split}\underset{\left(\beta,b,\xi\right)}{\min}&\;\;\;\;\frac{1}{2}\|\beta\|^{2}+K\sum_{i=1}^{n}\sum_{j=1}^{T_{i}}\xi_{i_{j}}+\lambda\sum_{i}^{n}g_{i}^{2}\\ \rm{s.t}&\;\;\;\;y_{i_{j}}(\beta^{\top}x_{i_{j}}+g_{i})\geq 1-\xi_{i_{j}},\text{ }i=1,\dots,n,\text{ }j=1,\dots,T_{i},\\ &\;\;\;\;\frac{1}{N}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{T_{i}}(s_{i_{j}}-\bar{s})(\beta^{\top}x_{i_{j}}+g_{i})\leq c,\\ \;&\;\;\;\;\frac{1}{N}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{T_{i}}(s_{i_{j}}-\bar{s})(\beta^{\top}x_{i_{j}}+g_{i})\geq-c,\\ &\;\;\;\;\xi_{i_{j}}\geq 0,\text{ }i=1,\dots,n,\text{ }j=1,\dots,T_{i},\end{split}$ (7)
with, $y_{i_{j}}$ being the label in the observation $j$ in group $i$ and $j\leavevmode\nobreak\ \in\leavevmode\nobreak\ [1,T_{i}]$, and $T_{i}$ the size of the group $i$. We refer to the problem above as Fair Mixed Effects Support Vector Machine (FMESVM) and without the fairness constraints as Mixed Effects Support Vector Machine (MESVM).
One alternative approach to incorporating random effects into grouped data analysis involves the use of dummy variables, also known as one-hot encoding as can be seen in [inbook234]. This strategy involves creating a separate binary variable for each group, taking a value of $1$ for observations within that group and 00 otherwise. While this method can effectively capture group-level variation, preliminary numerical tests have demonstrated that our proposed approach (7) offers significant advantages in terms of both computational efficiency and memory usage, making it a more resource-friendly option for large datasets and complex models in both situations where the datasets present unfairness and when they do not.
The following preliminar numerical experiment employed a dataset consisting of $100000$ points with the training set having 3 to 5 points per group and systematically varying the number of groups within the data, considering configurations with $2$, $10$, $50$, $500$, $1000$, $1250$, $2000$, $2500$, $3125$, $4000$, and $5000$ groups.
The Figures 1 and 2 shows two different analyses of the data. The top line visualizes a second-order polynomial fit. The bottom line, on the other hand, creates a Performance Profile proposed by [DolanMore02] of both approaches.
Figure 1. Memory comparison in Bytes.
Figure 2. Time comparison in Microsecond.
The next section dives into numerical tests, demonstrating the efficacy of the proposed methodology.
4. Simulation Study
First we present the step-by-step strategy used to create the datasets and to conduct the numerical experiments. Using the Julia 1.9 ([BezansonEdelmanKarpinskiShah17]) language with the packages Distributions ([distributions]), DataFrames ([DataFrames]), MLJ ([MLJ]) and MKL ([MKL]), we generate the following parameters:
  • Number of points: Number of points in the dataset;
  • $\beta^{\prime}s$: The fixed effects;
  • $g^{\prime}s$: The random effects with distribution $N(0,2)$;
  • Data points: The covariate vector associated with fixed effects with distribution $N(0,1)$;
  • $c$: Threshold from Fair Support Vector Machine;
  • seed: Random seed used in the generation of data;
  • Train-Test split: Approximately 0.4% of the dataset was used for the training set, and 99.6% for the test set. This percentage was due to the fact that we randomly selected 3 to 5 points from each cluster for the training set.
then the predictions of the synthetic dataset were computed using
$m=\beta^{T}x+g$ (8)
in tests where the dataset has random effects, and
$m=\beta^{T}x$ (9)
in the tests where the dataset has just fixed effects. Since $m\in[-\infty,\infty]$, we project the value to $[0,1]$ using the function:
$M=\frac{exp(m)}{1+exp(m)}.$ (10)
Finally,
$y=Binomial(1,M).$ (11)
Thus, our datasets are ready.
Regarding the tests, the comparisons will be made between the following optimizations problems:
  1. (1)
    Mixed Effects Suport Vector Machine (MESVM);
  2. (2)
    Fair Mixed Effects Suport Vector Machine (FMESVM);
  3. (3)
    Suport Vector Machine (SVM);
  4. (4)
    Fair Suport Vector Machine (SVMF).
comparing the accuracy and the disparate impact between them. The test were conducted on a computer with an Intel Core i9-13900HX processor with a clock speed of 5.40 GHz, 64 GB of RAM, and Windows 11 operating system, with 64-bit architecture.
To compute accuracy, first we need compute the predictions so we use equations (8) - (11) with,
$\hat{y}=\begin{cases}1\hskip 11.38092pt\text{if }\hskip 1.42271ptm\geq 0.5\\ -1\hskip 11.38092pt\text{if }\hskip 1.42271ptm<0.5\end{cases}\,,$
and given the true label of all points, we can distinguish them into four categories: true positive (TP) or true negative (TN) if the point is classified correctly in the positive or negative class, respectively, and false positive (FP) or false negative (FN) if the point is misclassified in the positive or negative class, respectively. Based on this, we can compute the accuracy, where a higher value indicates a better classification, as follows,
$AC=\dfrac{TP+TN}{TP+TN+FP+FN}\in[0,1].$
To compute the Disparate Impact of a specific sensitive feature $s$ we use the following equation as can be seen in [DIII],
$DI=\min\Big{(}\dfrac{p(\hat{y}|s=1)}{p(\hat{y}|s=0)},\dfrac{p(\hat{y}|s=0)}{p(\hat{y}|s=1)}\Big{)}\in[0,1].$
Disparate impact, as a measure, should be equal to 1. This indicates that discrimination does not exist. Values greater or lower than 1 suggest that unwanted discrimination exists. However, $DI=2$ and $DI=0.5$ represent the same level of discrimination, although in the first case, the difference between the perfect value is 1, and in the latter case, it is 0.5. To avoid such situations, we use the minimum of two possible values.
The parameters for the Unfair cases are:
  • $\beta$’s $=[-1.5,0.4,0.8,0.5,1.5]$;
  • $g$’s: 100 clusters with $b_{i}\sim N(0,Q)$, with $i\in[1,100]$;
  • $c=10^{-3}$;
  • $K=1$;
  • $\lambda=1$.
And for the fair cases are:
  • $\beta$’s $=[-1,1,2,1,0.1]$;
  • $g$’s: 100 clusters with $b_{i}\sim N(0,Q)$, with $i\in[1,100]$;
  • $c=10^{-3}$;
  • $K=1$;
  • $\lambda=1$.
with $Q$ being $2$ for cases with random effect and 00 for cases without random effect.
The following figures present experimental results on four different datasets with 1000 samples each. The datasets are presented individually, and the changes between samples occur only in the training set. Each figure contains four corresponding images representing accuracy and disparate impact for all possible datas. All figures were created using the Plots and PlotlyJS packages, developed by [Plots].
Figure 3. Accuracy.
Figure 4. Disparate Impact.
Figure 3 demonstrates that our proposed approach consistently outperformed alternative methods in scenarios with random effects. In settings without random effects, both methods achieved comparable accuracy, as anticipated. However, a slight decrease in accuracy was observed when introducing unfairness in the data. This is understandable, as our approach must balance addressing both random effects and unfairness simultaneously.
Figure 4 demonstrates that our approach consistently yields improved disparate impact metrics when applied to datasets containing inherent biases. Conversely, in scenarios where the underlying data exhibits no inherent bias, our approach produces equivalent outcomes, as there is no inherent disparity to mitigate.
5. Aplications
In this set of experiments, we will test the Adult dataset. To test the method’s efficiency, we created groups based on individuals age and marital status, and a sensitive feature, gender ([speicher2018unified]). The Adult dataset is a famous tool in machine learning, where the goal is predicting whether the individuals earn more ($y=1$) or less ($y=0$) than $50000$ USD annually. Were conducted 1000 samples with 0.5% of the data as the training set and the remaining data as the test set.
The features used in the prediction process are the follows:
  • Age: The age of the individual in years;
  • Capital_gain: Capital gain in the previous year;
  • Capital_loss: Capital loss in the previous year;
  • Education: Highest level of education achieved by the individual;
  • Education_num: A numeric form of the highest level of education achieved;
  • Fnlwgt: An estimate of the number of individuals in the population with the same demographics as this individual;
  • Hours_per_week: Hours worked per week;
  • Marital_status: The marital status of the individual;
  • Native_country: The native country of the individual;
  • Occupation: The occupation of the individual;
  • Race: The individual’s race;
  • Relationship: The individual’s relationship status;
  • Gender: The individual’s gender;
  • Workclass: The sector that the individual works in.
Figure 5. Accuracy.
Figure 6. Disparate Impact.
As can be seen in Figure 5, in this set of experiments, we obtained a better accuracy in MESVM and FMESVM in comparison to regular SVM since this one not account for random effects.
Figure 6 show that we also obtained an improvement in the disparate impact on the Fair algorithms. Note that make sense since we have a unfair population.
Upon examining all the tests, we were able to observe an improvement in Disparate Impact in ($100\%$) of the cases.
6. Conclusion
This study investigated a novel approach to mitigating disparate impact, a fairness issue in machine learning models, while simultaneously addressing mixed effects. We introduce a novel Fair Mixed Effects Support Vector Machine (FMESVM) algorithm that tackles both concerns cohesively, overcoming the limitations of existing methods often dedicated to separate problem solving. This integrated approach tailors treatments to the specific demands of each issue, ensuring optimal performance.
Employing the widely respected Support Vector Machine (SVM) for binary classification, the FMESVM framework incorporates mixed effects within the SVM setting and deploys novel regularization techniques to effectively mitigate disparate impact. Extensive evaluation across diverse datasets and metrics demonstrates the success of our proposed method in demonstrably reducing disparate impact while maintaining or minimally compromising overall accuracy.
For comprehensive experimentation, we systematically explored all possible scenarios involving the two concerns: datasets exhibiting both, only unfairness with random effects, only fairness with random effects, and neither issue. This approach yielded expected results, with each combination directly impacting accuracy or disparate impact as predicted.
The FMESVM presents a significant advancement in fairness-aware machine learning by comprehensively addressing disparate impact and mixed effects through a unified framework. This paves the way for the development of more robust and ethical machine learning models with broader applicability.
Acknowledgements
The authors are grateful for the support of the German Federal Ministry of Education and Research (BMBF) for this research project, as well as for the “OptimAgent Project”.
We would also like to express our sincere appreciation for the generous support provided by the German Research Foundation (DFG) within Research Training Group 2126 “Algorithmic Optimization”.
[…]